# Probing Dark Matter Substructures with Free-Form Modelling: A Case Study of the ‘Jackpot’ Strong Lens

### v1

@@ 4.2.1 Initialisation with True Lens and Source Models
Figure 6: Results of simultaneous source and perturbative lensing potential modelling to directly fit the observed lensing image, using the input main lens mass and source light models as the starting point. From left to right: the first row displays the data ($\boldsymbol{d}$), model ($\boldsymbol{d}^{m}$), and normalised residual (($\boldsymbol{d}-\boldsymbol{d}^{m})/\boldsymbol{n}$) images. The second row shows the reconstructed source ($\boldsymbol{s}^{\mathrm{M}}$), the lensing potential correction ($\delta\boldsymbol{\psi}^{M}$), and the corresponding convergence correction ($\delta\boldsymbol{\kappa}^{\mathrm{M}}$).
@@ 4.3 Characterising the Subhalo in the Jackpot Lens via a Free-Form Approach
To apply the potential correction method to the Jackpot lens, we first generate an annular mask around the lensed arc by connecting pixels with a lensed image SNR greater than 3 and expanding the connected region by 5 pixels. Next, we use the best-fit mass model derived by He et al. (2024), where the main lens’s mass is described by the EPL plus shear model, as the starting point to rerun a macro model that does not account for any perturbations. This refined macro model employs a pixelized source model with Kmeans pixelization and AdaptiveBrightnessSplit regularisation. The best-fit mass model and the corresponding pixelized source reconstruction are then used to construct the $\mathbf{L}\left(\psi_{p},s_{p}\right)$ matrix and solve for the unknown lensing potential perturbations, as detailed in Section 2.3. The hyperparameters associated with the source model, including Kmeans pixelization and AdaptiveBrightnessSplit regularisation, are fixed to the best-fit values obtained from the macro model when performing the potential correction. Allowing these hyperparameters to vary as free parameters is numerically unstable and renders the matrix equation 21 unsolvable.
Figure 10 presents the potential correction results for the Jackpot lens. Our model fits the observed lensing image exceptionally well, reducing the remaining image residuals to the noise level. A localised mass perturbation attributed to a subhalo is evident in the lensing potential perturbation map (manifested as a negative lensing potential well) and the corresponding convergence perturbation map (manifested as a positive mass clump). The central position of this localised mass perturber, defined as the pixel position with the maximum value in the convergence perturbation map, is $(y,x)=(-0.65,-1.05)$, consistent with the detection by Vegetti et al. (2010).
[…]
@@
While the flexible nature of the free-form model allows it to avoid many systematics present in parametric approaches, this flexibility can also make the performance of free-form models less stable than parametric ones. Therefore, we conduct several tests to evaluate the robustness of our potential correction results for the Jackpot lens. First, we applied a larger annular mask to the Jackpot lens and performed the potential correction again, finding that the properties of the reconstructed subhalo remained unchanged. Second, the potential correction results reveal additional features in addition to the detected subhalo. For instance, a dipole-like structure appears near the detected subhalo (in the top-right region), and a negative convergence peak is observed at $(x,y)=(-1.0,-1.0)$. To investigate the origin of these features, we employ a parametric approach analogous to that used by Despali et al. (2024) to model the Jackpot lens and simulate Jackpot-like mock lenses based on this model. Applying our potential correction method to these mock lenses reproduces the same additional features. Further analysis suggests that these arise from initiating the potential correction with a biased estimate of the main lens mass and source light provided by the macro model. Third, we simulate two Jackpot-like mock lenses with subhalo properties given by Despali et al. (2024) and Nightingale et al. (2024), respectively. After applying our potential correction method to these mock lenses, we observe a localised perturber in both cases. Moreover, the perturber exhibits a more compact structure in the mock lens generated using the model of Despali et al. (2024). This test demonstrates our potential correction method’s ability to reveal the compactness of a localised perturber. A detailed description of the robustness tests conducted here is provided in Appendix B.
@@
Figure 11: Convergence maps of the SLAC0946+1006 subhalo from different studies. Left: A $\sim 10^{10}\,M_{\odot}$ NFW subhalo with a high concentration of $\sim 200$, centered at $(-1.040,-0.651)$ (Despali et al., 2024). Middle: A $\sim 10^{11.5}\,M_{\odot}$ NFW subhalo following the standard mass-concentration relation, located at $(-1.00,-1.10)$ (Nightingale et al., 2024). Right: The perturbative convergence map resulting from our potential correction. Black contours indicate the lensed image morphology.
@@ 5.1 Enhanced Regularisation Approaches for Lensing Potential Perturbations
Our potential correction framework is based on the quadratic form of the regularisation function, which allows the unknown perturbations in the lensing potential to be solved linearly. It has been demonstrated that quadratic regularisation with the Matérn kernel can effectively recover various types of lensing potential perturbations, whether extended or localised. Despite these impressive results, we observe certain artefacts in the potential correction results, such as random negative-convergence features, which have also been noted in previous works (Vegetti et al., 2010). These artefacts stem from the inherent nature of quadratic regularisation (or L2 regularisation), which promotes smoothness by shrinking all feature values (specifically, the lensing potential perturbations at each pixel) without setting any to zero. To improve the potential correction method, an alternative approach to imposing smoothness priors may involve setting the lensing potential perturbation values at each pixel to zero, unless the data indicate otherwise. This requirement aligns with the purpose of L1 regularisation. However, utilising L1 regularisation disrupts the linear inversion framework, preventing us from using equation 16 to solve for the lensing potential perturbation. Instead, we must determine the lensing potential perturbation values by directly optimising the penalty function illustrated in equation 15. This approach introduces a substantial number of nonlinear modelling parameters (several thousand), which are challenging for optimisers that do not leverage gradient information to navigate parameter space. Since our current code lacks support for auto-differentiation, we are unable to compute the gradient of the likelihood with respect to the modelling parameters, rendering our code incompatible with L1 regularisation at this time. Several open-source lens modelling software packages have incorporated auto-differentiation (Gu et al., 2022; Galan et al., 2022), and PyAutoLens plans to integrate this functionality soon. We plan to incorporate L1 regularisation (and potentially more general forms) into our code in future work..
@@ 5.2 Quantifying Uncertainties in Lensing Potential Perturbations
The artefacts in our current potential correction results are unavoidable due to the limitations of quadratic regularisation schemes. However, if we can quantify the uncertainties in the model-derived lensing potential perturbations and demonstrate that the signal in these artefact regions is highly uncertain, these artefacts will not impede the interpretation of the potential correction results. Given a set of hyperparameters for the source and lensing potential perturbation regularisations, which are best determined by maximizing the Bayesian evidence, we can estimate the uncertainties of the pixelized source and lensing potential perturbations using the covariance matrix, the mathematical form of which is provided in Equation 17777Replace $\mathbf{L}_{\delta\psi}$ with $\mathbf{L}$.. This covariance matrix contains numerous non-zero off-diagonal elements, indicating a high degree of correlation between the pixelized source and lensing potential perturbation reconstructions. Since our primary interest lies in determining whether a mass perturbation structure is genuine, we aim to quantify the uncertainties of pixelized mass perturbations and define an SNR-like quantity to evaluate the fidelity of the perturbation signal. However, we currently lack a robust method to translate the covariance matrix of the lensing potential perturbations into uncertainties for the mass perturbations. Consequently, we resort to Monte Carlo simulations to estimate the uncertainties of pixelized mass perturbations, as detailed in Appendix B.
@@ 5.3 Incorporating Lens Light Modelling within Potential Correction Algorithms
Here, the matrix $\mathbf{L_{l}}(\xi_{l})$ has dimensions $[N_{d},N_{l}]$, with each column representing the model image formed by each Gaussian model with a "unit" normalization value; $\xi_{l}$ denotes the nonlinear parameters corresponding to the centre and ellipticities of the Gaussian models in this instance. The column vector $\boldsymbol{l}$ contains $N_{l}$ elements, representing the normalisation factors for each Gaussian model. Equation 21 can now be extended to include lens light modelling as follows:
@@
where $\mathbf{I}_{l}$ is an identity matrix of dimensions $[N_{l},N_{l}]$, and $\epsilon$ is a small constant (e.g., $10^{-5}$). Equations 21 and 39 retain the same mathematical form; thus, the linear inversion framework outlined in Section 2.3 remains applicable for potential corrections that incorporate lens light modelling.
@@ 5.4 Concurrent Modelling of Main Lens and Lensing Potential Perturbations
In Section 4.2.2, we demonstrated that the perturbations derived using the potential correction method do not generally match the input ground truth, as the macro model can vary to absorb some of the mass perturbations. To enable the potential correction method to recover the underlying true perturbations, a natural approach would be to simultaneously fit the main lens mass and the lensing potential perturbations (e.g. Vernardos & Koopmans, 2022). Although we cannot currently perform this simultaneous fitting due to a fixed main lens mass model being hard-coded in our programme, we observe that starting the potential correction from the input true main lens mass model yields higher Bayesian evidence than starting from the macro model’s estimate. This suggests that simultaneous fitting of the main lens mass and lensing potential perturbations might be feasible. We plan to explore this possibility in future work.
@@ Acknowledgements
This work was supported by the National Key R&D Program of China (grant number 2022YFF0503403), the National Natural Science Foundation of China (No. 11988101), the K.C.Wong Education Foundation, and the science research grants from China Manned Space Project with No.CMS-CSST-2021-B01. XYC acknowledges the support of the National Natural Science Foundation of China (No. 12303006). RL is also supported by the National Natural Science Foundation of China (Nos 11773032, 12022306). This work was supported in the UK by STFC via grant ST/X001075/1, the UK Space Agency via grant ST/W002612/1, and the European Research Council via grant DMIDAS (GA 786910). We express our gratitude to ChatGPT, an AI model developed by OpenAI, for its assistance in polishing the English of this paper.
@@
Appendix A Optimal Hyperparameter Values for the Matern Kernel
For the interested reader, we present the optimal hyperparameter values used to regularise lensing potential perturbations with the Matérn kernel in Table 4. These values were determined objectively by sampling the Bayesian evidence.

### v2

@@ 4.2.1 Initialisation with True Lens and Source Models
Figure 6: Results of simultaneous source and perturbative lensing potential modelling to directly fit the observed lensing image, using the input main lens mass and source light models as the starting point. From left to right: the first row displays the data ($\boldsymbol{d}$), model ($\boldsymbol{d}^{m}$), and normalised residual $($$\boldsymbol{d}-\boldsymbol{d}^{m})/\boldsymbol{n}$) images. The second row shows the reconstructed source ($\boldsymbol{s}^{\mathrm{M}}$), the lensing potential correction ($\delta\boldsymbol{\psi}^{M}$), and the corresponding convergence correction ($\delta\boldsymbol{\kappa}^{\mathrm{M}}$). The ‘plus’ symbol in the top-left and top-middle panels marks the position of the input subhalo.
@@ 4.3 Characterising the Subhalo in the Jackpot Lens via a Free-Form Approach
To apply the potential correction method to the Jackpot lens, we first generate an annular mask around the lensed arc by connecting pixels with a lensed image SNR greater than 3 and expanding the connected region by 5 pixels. Next, we use the inferred mass model derived by He et al. (2024), where the main lens’s mass is described by the EPL plus shear model, as the starting point to rerun a macro model that does not account for any perturbations. This refined macro model employs a pixelized source model with Kmeans pixelization and AdaptiveBrightnessSplit regularisation. The inferred mass model and the corresponding pixelized source reconstruction are then used to construct the $\mathbf{L}\left(\psi_{p},s_{p}\right)$ matrix and solve for the unknown lensing potential perturbations, as detailed in Section 2.3. The hyperparameters associated with the source model, including Kmeans pixelization and AdaptiveBrightnessSplit regularisation, are fixed to the values obtained from the macro model when performing the potential correction. Allowing these hyperparameters to vary as free parameters is numerically unstable and renders the matrix equation 12 unsolvable.
Figure 10 presents the potential correction results for the Jackpot lens. Our model fits the observed lensing image exceptionally well, reducing the remaining image residuals to the noise level. A localised mass perturbation attributed to a subhalo is evident in the lensing potential perturbation map (manifested as a negative lensing potential well) and the corresponding convergence perturbation map (manifested as a positive mass clump). The central position of this localised mass perturber, defined as the pixel position with the maximum value in the convergence perturbation map, is (y,x)=(0.65,1.05)(y,x)=(-0.65,-1.05), consistent with the detection by Vegetti et al. (2010).
Minor et al. (2017) define a model-independent method to measure the “robust mass” of the subhalo, which is the aperture mass within the so-called robust radius $R_{\text{robust}}$. This radius can be calculated based on the geometric perturbations induced by the subhalo on the main lens’s critical lines. For the subhalo in the Jackpot lens, $R_{\text{robust}}\approx 1$ kpc (Minor et al., 2021), as indicated by the black circle in Figure 10. This yields a robust mass of $3\times 10^{8}M_{\odot}$ based on our potential correction result. However, this value is an order of magnitude lower than the result reported by Minor et al. (2021), who explicitly used a truncated NFW model to describe the subhalo’s mass, deriving a robust mass of $\sim 2-3.7\times 10^{9}M_{\odot}$ ($\sim 95\%$ confidence level). We attribute this discrepancy to the limitations of our regularisation scheme, which, although it clearly performs better than the alternatives (as shown in Figure 4), still lacks the dynamic range needed to reconstruct highly compact perturbers. Specifically, we impose a constant regularisation strength for lensing potential perturbations across the entire modelling region. Although Bayesian evidence determines a relatively strong regularisation strength to suppress noise-like features in regions without significant perturbations, this approach simultaneously oversmooths the structure of highly compact perturbers. Notably, this limitation is analogous to the challenges faced when modelling clumpy sources with steep brightness gradients. In Section 5.1, we discuss how adaptive regularisation for lensing potential perturbations could mitigate these systematics.
@@
While the flexible nature of the free-form model allows it to avoid many systematics present in parametric approaches, this flexibility can also make the performance of free-form models less stable than parametric ones. Therefore, we conduct several tests to evaluate the robustness of our potential correction results for the Jackpot lens. First, we applied a larger annular mask to the Jackpot lens and performed the potential correction again, finding that the properties of the reconstructed subhalo remained unchanged. Second, the potential correction results reveal additional features in addition to the detected subhalo. For instance, a dipole-like structure appears near the detected subhalo (in the top-right region), and a negative convergence peak is observed near $(x,y)\approx(1.0,1.0)$. To investigate the origin of these features, we employ a parametric approach analogous to that used by Despali et al. (2024) to model the Jackpot lens and simulate Jackpot-like mock lenses based on this model. Applying our potential correction method to these mock lenses reproduces the same additional features. Further analysis suggests that these arise from initiating the potential correction with a biased estimate of the main lens mass and source light provided by the macro model. Third, we simulate two Jackpot-like mock lenses with subhalo properties given by Despali et al. (2024) and Nightingale et al. (2024), respectively. After applying our potential correction method to these mock lenses, we observe a localised perturber in both cases. Moreover, the perturber exhibits a more compact structure in the mock lens generated using the model of Despali et al. (2024). This test demonstrates our potential correction method’s ability to reveal the compactness of a localised perturber. A detailed description of the robustness tests conducted here is provided in Appendix C.
@@
Figure 11: Convergence maps of the SLAC0946+1006 subhalo from different studies. Left: A $\sim 10^{10}\,M_{\odot}$ NFW subhalo with a high concentration of $\sim 200$, centered at (1.040,0.651)(-1.040,-0.651) (Despali et al., 2024). Middle: A $\sim 10^{11.5}\,M_{\odot}$ NFW subhalo following the standard mass-concentration relation, located at (1.00,1.10)(-1.00,-1.10) (Nightingale et al., 2024). Right: The perturbative convergence map resulting from our potential correction. Black contours indicate the lensed image morphology.
@@ 5.1 Enhanced Regularisation Approaches for Lensing Potential Perturbations
Our potential correction framework is based on the quadratic form of the regularisation function, which allows the unknown perturbations in the lensing potential to be solved linearly. It has been demonstrated that quadratic regularisation with the Matérn kernel can effectively recover various types of lensing potential perturbations, whether extended or localised. Despite these impressive results, we observe certain artefacts in the potential correction results, such as random negative-convergence features, which have also been noted in previous works (Vegetti et al., 2010). These artefacts stem from the inherent nature of quadratic regularisation (also known as L2 regularisation or ridge regression), which promotes smoothness by shrinking all feature values (specifically, the lensing potential perturbations at each pixel) without setting any to zero (Hoerl and Kennard, 1970b, a). To improve the potential correction method, an alternative approach to imposing smoothness priors may involve setting the lensing potential perturbation values at each pixel to zero, unless the data indicate otherwise. This requirement aligns with the purpose of L1 regularisation (or LASSO; Tibshirani, 1996). However, utilising L1 regularisation disrupts the linear inversion framework, preventing us from using equation 9 to solve for the lensing potential perturbation. Instead, we must determine the lensing potential perturbation values by directly optimising the penalty function illustrated in equation 40. This approach introduces a substantial number of nonlinear modelling parameters (several thousand), which are challenging for optimisers that do not leverage gradient information to navigate parameter space. Since our current code lacks support for auto-differentiation, we are unable to compute the gradient of the likelihood with respect to the modelling parameters, rendering our code incompatible with L1 regularisation at this time. Several open-source lens modelling software packages have incorporated auto-differentiation (Gu et al., 2022; Galan et al., 2022; Cao et al., 2025), and PyAutoLens plans to integrate this functionality soon. We plan to incorporate L1 regularisation (and potentially more general forms) into our code in future work.
@@ 5.2 Quantifying Uncertainties in Lensing Potential Perturbations
The artefacts in our current potential correction results are unavoidable due to the limitations of quadratic regularisation schemes. However, if we can quantify the uncertainties in the model-derived lensing potential perturbations and demonstrate that the signal in these artefact regions is highly uncertain, these artefacts will not impede the interpretation of the potential correction results. Given a set of hyperparameters for the source and lensing potential perturbation regularisations, which are best determined by maximizing the Bayesian evidence, we can estimate the uncertainties of the pixelized source and lensing potential perturbations using the covariance matrix, the mathematical form of which is provided in Equation 41777Replace $\mathbf{L}_{\delta\psi}$ with $\mathbf{L}$.. This covariance matrix contains numerous non-zero off-diagonal elements, indicating a high degree of correlation between the pixelized source and lensing potential perturbation reconstructions. Since our primary interest lies in determining whether a mass perturbation structure is genuine, we aim to quantify the uncertainties of pixelized mass perturbations and define an SNR-like quantity to evaluate the fidelity of the perturbation signal. However, we currently lack a robust method to translate the covariance matrix of the lensing potential perturbations into uncertainties for the mass perturbations. Consequently, we resort to Monte Carlo simulations to estimate the uncertainties of pixelized mass perturbations, as detailed in Appendix C.
@@ 5.3 Incorporating Lens Light Modelling within Potential Correction Algorithms
Here, the matrix $\mathbf{L_{l}}(\xi_{l})$ has dimensions $N_{d},N_{l}]$, with each column representing the model image formed by each Gaussian model with a "unit" normalization value; $\xi_{l}$ denotes the nonlinear parameters corresponding to the centre and ellipticities of the Gaussian models in this instance. The column vector $\boldsymbol{l}$ contains $N_{l}$ elements, representing the normalisation factors for each Gaussian model. Equation 12 can now be extended to include lens light modelling as follows:
@@
where $\mathbf{I}_{l}$ is an identity matrix of dimensions $N_{l},N_{l}]$, and $\epsilon$ is a small constant (e.g., $10^{-5}$). Equations 12 and 30 retain the same mathematical form; thus, the linear inversion framework outlined in Section 2.3 remains applicable for potential corrections that incorporate lens light modelling.
@@ 5.4 Concurrent Modelling of Main Lens and Lensing Potential Perturbations
In Section 4.2.2, we demonstrated that the perturbations derived using the potential correction method do not generally match the input ground truth, as the macro model can vary to absorb some of the mass perturbations. To enable the potential correction method to recover the underlying true perturbations, a natural approach would be to simultaneously fit the main lens mass and the lensing potential perturbations (e.g. Vernardos and Koopmans, 2022). Although we cannot currently perform this simultaneous fitting due to a fixed main lens mass model being hard-coded in our programme, we observe that starting the potential correction from the input true main lens mass model yields higher Bayesian evidence than starting from the macro model’s estimate. This suggests that simultaneous fitting of the main lens mass and lensing potential perturbations might be feasible. We plan to explore this possibility in future work.
@@ Acknowledgements
We thank the referee for the helpful comments and suggestions that have improved this paper. This work was supported by the National Key R&D Program of China (grant number 2022YFF0503403), the National Natural Science Foundation of China (No. 11988101), the K.C.Wong Education Foundation, the Ministry of Science and Technology of China (No. 2020SKA0110100), and the science research grants from China Manned Space Project with Nos.CMS-CSST-2021-B01 and CMS-CSST-2025-A03. XYC acknowledges the support of the National Natural Science Foundation of China (No. 12303006). RL is also supported by the National Natural Science Foundation of China (Nos 11773032, 12022306), the CAS Project for Young Scientists in Basic Research (No. YSBR-062). This work was supported in the UK by STFC via grant ST/X001075/1, the UK Space Agency via grant ST/W002612/1, and the European Research Council via grant DMIDAS (GA 786910). We express our gratitude to ChatGPT, an AI model developed by OpenAI, for its assistance in polishing the English of this paper.
@@
Appendix A Bayesian Framework for Potential Correction
Equation 6 shows that $\delta\boldsymbol{d}$ and $\boldsymbol{\delta\psi}$ are connected via a linear mapping. The most likely solution for $\boldsymbol{\delta\psi}$ can be derived by maximizing the likelihood
where
$=\frac{1}{2}\chi^{2},$
and
The most likely solution $\boldsymbol{\delta\psi}_{\mathrm{ml}}$ is usually degraded by noise and is ill-posed (Warren and Dye, 2003). An additional prior $P(\delta\boldsymbol{\psi}\mid\boldsymbol{g}_{\delta\psi},\boldsymbol{\xi}_{\delta\psi})$ is needed to impose smoothness on $\delta\boldsymbol{\psi}$, a process commonly called “regularisation”. Here, $\boldsymbol{g}_{\delta\psi}$ represents the type of regularisation, and the associated model parameters are denoted by the vector $\boldsymbol{\xi}_{\delta\psi}$. The prior can be expressed as:
where $Z_{\delta\psi}(\boldsymbol{\xi}_{\delta\psi})=\int\mathrm{d}^{N_{p}}(\delta\boldsymbol{\psi})\exp(-E_{\mathrm{\delta\psi}})$ is the normalization factor of the prior probability density function. The function $E_{\delta\psi}$, also known as the regularisation function, typically takes a quadratic form: $E_{\delta\psi}(\delta\boldsymbol{\psi})=\frac{1}{2}\delta\boldsymbol{\psi}^{\mathrm{T}}\mathbf{R_{\delta\psi}}\delta\boldsymbol{\psi}$, where $\mathbf{R_{\delta\psi}}$ represents the Hessian of $E_{\delta\psi}$. This form allows $\delta\boldsymbol{\psi}$ to be directly solved through matrix inversion (Warren and Dye, 2003). For more detailed information on the specific form of the regularisation function, refer to Section 2.4.
This yields Equation 9:
$\boldsymbol{\delta\psi}_{\mathrm{MAP}}\;=\;\left(\mathbf{L}^{\mathrm{T}}_{\delta\psi}\,\mathbf{C}_{\mathrm{D}}^{-1}\,\mathbf{L}_{\delta\psi}+\mathbf{R}_{\delta\psi}\right)^{-1}\mathbf{L}^{\mathrm{T}}_{\delta\psi}\,\mathbf{C}_{\mathrm{D}}^{-1}\,\delta\boldsymbol{d},$
which indicates that, given the macro-model outputs $\{\delta\boldsymbol{d},\,\mathbf{L}_{\delta\psi}\}$ and an arbitrary choice of regularisation $\mathbf{R}_{\delta\psi}(\boldsymbol{\xi}_{\delta\psi},\boldsymbol{g}_{\delta\psi})$, the most probable solution, $\boldsymbol{\delta\psi}_{\mathrm{MAP}}$, can be obtained via linear matrix inversion. The covariance matrix ($\Sigma_{\delta\psi}$), which describes the uncertainty of the most probable solution ($\boldsymbol{\delta\psi}_{\mathrm{mp}}$), is derived by Suyu et al. (2006) and follows:
Here, the evidence term $P(\delta\boldsymbol{d}\mid\mathbf{L}_{\delta\psi},\boldsymbol{g}_{\delta\psi})$ marginalizes over the hyperparameters of the regularisation and can be used to rank different model choices of $\{\mathbf{L}_{\delta\psi},\boldsymbol{g}_{\delta\psi}\}$. Since a non-informative prior is usually taken for $P(\boldsymbol{\xi}_{\delta\psi})$, finding the most probable solution for $\boldsymbol{\xi}_{\delta\psi}$ is equivalent to maximizing $P(\delta\boldsymbol{d}\mid\mathbf{L_{\delta\psi}},\boldsymbol{\xi}_{\delta\psi},\boldsymbol{g}_{\delta\psi})\equiv\mathcal{E}$, whose explicit form is given in Equation 10.
Appendix B Optimal Hyperparameter Values for the Matern Kernel
In this work, the Matérn kernel is employed to regularise the pixelised potential corrections. As defined in Equation 19, the Matérn kernel is characterised by three hyperparameters: a coefficient $\lambda^{\delta\psi}$ that controls the overall smoothing strength; a characteristic length-scale $\rho$ (in units of arcseconds) that governs the smoothing scale; and a parameter $\nu$ that determines the order of the smoothing, ensuring the reconstructed field is $\left\lceil\nu\right\rceil-1$ times differentiable. We assign non-informative priors to these three hyperparameters: $\lambda^{\delta\psi}\sim\mathcal{L}(10^{-6},10^{6})$, $\rho\sim\mathcal{L}(10^{-4},10^{3})$, and $\nu\sim\mathcal{U}(0.5,10.0)$. The optimal values for these hyperparameters, presented in Table 4, were determined by sampling the Bayesian evidence.